OpenStudy (horsegirl27):

Please help! So, I know how to find the slope when given 2 sets of points. I can also figure out the slope when looking at a graph. But for problems such as 16x+8y=192, how do I find the slope??

OpenStudy (anonymous):

solve for y. Once you have it in y=mx+b form the slope is m

OpenStudy (horsegirl27):

Okay, thanks. So I'm trying to solve for y and put it in the other form?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

rearrange the equation so it looks like y=mx+b ... you may need to do some division because we need y by itself.

OpenStudy (horsegirl27):

What should I get when I do that? (this isn't a homework problem, I'm just finding practice problems to get more practice) I think if I see all the steps I may understand.

OpenStudy (usukidoll):

given: 16x+8y=192, we know that y = mx+b is the slope-intercept form so we need to rearrange 16x+8y=192 to make it look like y =mx+b so since the y portion needs to stay on the left what do we need to do to 16x?

OpenStudy (horsegirl27):

Divide, right?

OpenStudy (usukidoll):

that's later on.

OpenStudy (usukidoll):

16x+8y=192 but we need it in the form of y=mx+b . The 8y stays on the left.. so what do we need to do 16x... we need to _______ 16x to both sides

OpenStudy (anonymous):